Grade 11General Physics

The sum of 2 forces at a point is 16N. If their resultant is normal to the smaller force and has a magnitude of 8N. Then the two forces are

Sanskriti Gangrade
8 Years agoGrade 11

1 Answer

Soumya Ranjan Mohanty
8 Years ago
Dear student,Assume the 2 forces be A & B, and the resultant be R.Given that, A + B = 16 NR^2 = B2 – A2 = (B – A)(B + A) = 16(B – A)(B - A) = R^2/16 = 4B – A = 16B + A = 4Solving,2B = 12B = 6NA = 10NSo the 2 forces are 6N and 10 N.[Ans]Regards,Soumya Ranjan Mohanty.